OpenStudy (tennistar):

Solve the proportion x over x squared minus 2 equals 2 over 3 times x plus 5.

OpenStudy (tennistar):



OpenStudy (anonymous):

like this? \[\huge \frac{x}{x^2-2}=\frac{2}{3x+5} \]

OpenStudy (tennistar):

yes

OpenStudy (anonymous):

ok... cross multiplying and you'll end up with: \[\huge x(3x+5)=2(x^2-2) \] \[\huge 3x^2+5x=2x^2-4 \] can you solve this quadratic equation from here?

OpenStudy (tennistar):

um, no

OpenStudy (anonymous):

ok... get this equation in standard form... that is make it look like this \(\large ax^2+bx+c=0 \)

OpenStudy (anonymous):

do some algebra to make the right side zero....

OpenStudy (tennistar):

what? Tennis terms?

OpenStudy (anonymous):

lol.... just kidding...

OpenStudy (tennistar):

oh

OpenStudy (anonymous):

your the tennis star..:)

OpenStudy (anonymous):

need help?

OpenStudy (anonymous):

whatever you do on one side, you do to the other... since we need to make the right side zero, subtract \(\large 2x^2-4 \) from both sides.. \[\huge 3x^2+5x=2x^2-4 \] \[\huge 3x^2+5x\color{red}{-(2x^2-4)}=2x^2-4 \color{red}{-(2x^2-4)}\] \[\huge 3x^2+5x\color{red}{-2x^2+4}=0\] \[\huge x^2+5x+4=0 \]